I have 46 pounds of apples, each one weighs 4 ounces, How many apples do I have?: I have 46 pounds of apples, each one weighs 4 ounces, How many apples do I have?
Answer by checkley77(3654)

(46*16)/4=736/4=184 apples.
Question 142250: I have 46 pounds of apples, each one weighs 4 ounces, How many apples do I have?: I have 46 pounds of apples, each one weighs 4 ounces, How many apples do I have?
Answer by rapaljer(2773)


Each pound equals 16 ounces. If the apples are 4 ounces each, then this is 4 apples per pound. Multiply the number of pounds times 4, and you should get 4*46 = 184 apples.

R^2